# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v4

@@ C.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample [109]. A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
@@
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic [110, 97]. When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
@@ C.3 Goodness-of-fit Test
Chi-square ($\chi^{2}$) test[111], Kolmogorov–Smirnov test (KS test) [112], and Anderson–Darling test [113] are some examples of the goodness-of-fit test [114]. Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well [115]. The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions [116].
@@ Appendix D Quantum Entropy for QRNG
Definition 3
The Shannon entropy [99, 100] of a discrete random variable $X$ is defined as
@@ Appendix E Randomness Testing Suites
To overcome this limitation, several randomness testing suites like NIST [65], Dieharder [66], AIS-$31$ [67], ENT [68] etc. have been proposed. Recently, Foreman et al. also suggested some test suite [117] for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation [2] to validate the randomness.

### v5

@@ C.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample casella2002statistical . A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
@@
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic ross2009introduction ; hogg2019introduction . When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
@@ C.3 Goodness-of-fit Test
Chi-square ($\chi^{2}$) test10.1214/aoms/1177729380 , Kolmogorov–Smirnov test (KS test) 16e7f618-c06b-3d10-8705-1086b218d827 , and Anderson–Darling test 10.1214/aoms/1177729437 are some examples of the goodness-of-fit test good . Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well 45e251e1-e12b-3bce-b1a2-da8f4c1c2382 . The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions 32ba378b-c6d3-36dc-9d3e-0b9814b089a4 .
@@ Appendix D Quantum Entropy for QRNG
Definition 3.
The Shannon entropy 6773024 ; 6773067 of a discrete random variable $X$ is defined as
@@ Appendix E Randomness Testing Suites
To overcome this limitation, several randomness testing suites like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. have been proposed. Recently, Foreman et al. also suggested some test suite foreman2024 for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation MonteCarlo to validate the randomness.